### v1

@@ 6.2.3 Low-redshift galaxies
The CSST main survey is expected to measure the morphology and structure properties of young globular clusters of different galaxy types and environments within hundreds of Mpc. The high-quality imaging capability of the CSST will provide an unprecedented large sample of galaxy nuclear star clusters (NSCs). The high-precision photometric capability will obtain the NSC color index, radius, mass, stellar population, star formation history, morphology, and complete distribution of NSC physical parameters. In this way, the correlation with the parameters of AGN dust ring can be established, the physical connection between the NSCs, bulges and massive black holes in a galaxy can be explored, and the physical process from the galaxy to the AGN transition region can be revealed.
@@
The outer regions of disk galaxies contain unique information about the physical processes that affect galactic evolution. Using the CSST high spatial resolution images, we can detect star clusters and single bright stars in the outer disk of galaxies. Considering the large sky coverage of CSST, the data at different radii of galaxies can be stacked, so that the areas with low star surface density can be detected. This can be helpful to answer two questions: the extension scale of galaxies and the physical mechanism that causes the growth of stellar disks. In addition, many galaxies have extended ultraviolet disks (XUV disks) at their outer edges [99]. CSST can be used to conduct UV-optical observations of a large number of XUV galaxies that can resolve single stars and study their physical properties.
@@
Star formation in galaxies is one of the main driving forces of galaxy growth and evolution, and is closely related to the mass structure, nuclear activity and other physical characteristics of galaxies. CSST’s multi-color imaging survey provides near-UV high-resolution deep-field images, which can effectively detect and analyze small-scale ($<1$ kpc) star formation and distribution in low-$z$ galaxies. CSST multi-color imaging survey is also expected to quantify the relationship between the star formation rate and molecular gas density on the sub-kpc scale, by exploring the group behavior and statistical properties of star formation in low-$z$ galaxies. Besides, CSST can conduct analyses on galaxies with strong radiation in the UV spectrum, which will help explain their origins [100].
@@ 6.2.4 Medium and high-redshift galaxies
When the resolution and S/N of galaxy images are low (such as when the galaxy is close to the detection limit), there are large errors in the measurement of its structural parameters. Therefore, it is necessary to use simulated images to test how the measurement errors of galaxy structural parameters vary with resolution and S/N. The mock galaxy images can be obtained by using galaxy image analysis code (e.g. GALFIT) or real multi-band galaxy images (e.g. HST CANDELS). By applying the measurement methods, we can obtain the trend of measurement deviations of different structural parameters of galaxies as the image resolution and S/N change, which will help to understand and correct the statistical analysis results of galaxy image structural parameters.
The morphology of a galaxy characterizes its internal structure and its evolutionary stage. Galaxies of different morphological types may have different formation and evolution processes. The CSST image survey will significantly increase the number of samples of medium- and high-$z$ galaxies, and study their formation and evolution history of the morphology. It is expected that the statistical distribution characteristics of galaxy morphology within the highest redshift of 2 can be constructed, and the overall evolution process of the Hubble sequence can be reproduced. Besides, galaxy morphological parameters have rich physical meanings, and they are related to the star formation history of galaxies, intergalactic interactions, and small-scale structures.
@@
CSST will study the formation and evolution of substructures of medium and high-$z$ galaxies. The multi-component decomposition and statistical analysis of medium- and high-$z$ galaxy images will be explored, and the deviation of multi-component decomposition can be evaluated by using CSST simulated image data. It is expect to obtain a multi-component decomposition catalog of a large sample of galaxy images, and conduct statistical analysis on physical parameters, such as structural parameters of bulges and disks, bulge mass fraction, and stellar population in galaxies of different masses and star formation rates, to reveal their formation and evolution history as a function of redshift. In addition, CSST will focus on the identification and measurement of bars and spiral arms in galaxies within $z=1$. By combining HST and JWST data, CSST can identify and measure the bar structure in the early stages of disk galaxies and reveal the role of substructures in galaxies in their mass growth and structural evolution. CSST will also study the physical properties and evolution of clumps within galaxies.
@@ 6.2.5 Galaxy evolution and assembly history
Using CSST multi-band high-resolution photometric and slitless spectroscopic surveys, combined with other multi-band survey data, we are able to measure important parameters, such as stellar mass, star formation rate, dust extinction, stellar population age, and metallicity, for a large number of galaxies at the low and medium redshifts. CSST can also measure the overall properties of a large number of complete galaxy samples, combined with galaxy pairs, merging galaxy statistics, and different local environment measurements, to construct the mass/luminosity function at different redshifts, as well as the scaling relationship between the mass-structure parameters-stellar population properties. Combined with numerical simulations or different types of galaxy evolution models, statistical modeling of these observations will greatly deepen our understanding of the galaxy stellar assembly history.
@@
Besides, since CSST can perform high-resolution wide-field survey, it will effectively probe the morphology and structure of galaxies and their dependence on the environment. CSST can carry out measurements of environments at different scales (including groups, filaments, voids, etc.), and study the impact of the environment on various properties of galaxies (including mass function, star formation rate, morphology, halo assembly history, gas, etc.), and their redshift evolution. CSST can also use its high-resolution photometric and spectroscopic survey to explore the formation and evolution history of galaxy groups and clusters, the formation and assembly history of the member galaxies, and the environmental effects of different redshifts and different types of galaxy groups/clusters on the formation and evolution of their member galaxies.
@@ 6.3.1 Stellar population
As a galaxy in the nearby Universe, the study of stellar populations in the Milky Way has made great progress in recent years. The release of the $Gaia$ astrometric catalog combined with ground-based stellar spectral survey data, such as the Large Sky Area Multi-Object Fiber Spectroscopic Telescope (LAMOST) and SDSS, provides important observational data for the study of the Milky Way. The study of stellar populations in the Milky Way is mainly carried out in star clusters, field stars, nuclear bulges, etc.
Some studies indicate that the Milky Way may be special in the properties of star populations than other galaxies, so the understanding of the formation and evolution history of galaxies cannot rely solely on the results of the Milky Way. The statistical results of star populations in different types of galaxies will be helpful to have a deeper understanding of the specific process of galaxy formation and evolution.
@@
Based on this sample, great progress will be made in the properties of various stellar populations in nearby galaxies. Important observational results of the variation of the initial mass function in different interstellar environments will be obtained. The evolution of stellar populations in different galaxies and the characteristics and evolution of stellar halos in different types of galaxies can be accurately measured. The number and types of satellite galaxies, and the number, types and spatial distribution of star clusters around different types of galaxies will be probed. The fundamentals of satellite galaxies and star clusters of different types of galaxies will be established, providing unprecedented complete data for the origin, merger and accretion process of satellite galaxies and constraints on the formation and evolution of star clusters.
@@ 6.3.2 Extinction and dust
The dust of various sizes, ranging from sub-nanometer to micron, is distributed in various environments from Solar System to the early Universe, and plays an important role in astrophysics research. Although interstellar dust accounts for only one thousandth of the baryonic mass of the Milky Way, it interacts with electromagnetic radiation in almost the entire frequency domain, thus comprehensively affecting all branches of astrophysics. The absorption and scattering of starlight by dust is collectively called interstellar extinction. The measurement of interstellar extinction and infrared re-radiation of the Milky Way and extragalactic galaxies is the key to restore the intrinsic brightness and SED of astronomical sources and determine the star formation rate on a galaxy scale.
@@
Using the CSST multi-band photometric and slitless spectroscopic observations of the Milky Way and nearby galaxies in multiple sky regions and interstellar sight lines, combined with the infrared and millimeter data and interstellar element abundances from the Infrared Astronomical Satellite (IRAS), Spitzer and $Planck$, interstellar dust models can be constructed. The size distribution, chemical composition and other characteristics of dust can be obtained. The relationship between dust properties and the interstellar physical and chemical environment will be analyzed, and the relationship between dust extinction and other important interstellar observational features (such as interstellar diffuse belts, aromatic infrared radiation spectra, X-ray “halos”, etc.) can be studied. The electronic absorption spectral characteristics of carbon dust in the CSST UV band, such as graphene, carbon nanotubes, nanodiamonds, fullerenes (such as C60), carbon chains and non-crystalline carbon, will be examined. In short, CSST will bring unprecedented opportunities for the study of extinction and dust physics.
@@ 6.4 Stars
CSST will obtain photometric data of billions of stars and hundreds of millions of stellar spectra during the ten-year survey. Its high spatial resolution ($0.074^{\prime\prime}$/pixel) and deep survey depth ($r\simeq 26$ mag) will enable us to observe single stars in the Milky Way and dozens or hundreds of nearby galaxies, such as the Andromeda Galaxy and the Triangulum Galaxy. By utilizing the unique advantages of CSST in the near-IR and near-UV bands and combining with other survey projects (such as LSST, HST, $Gaia$, etc.), we can obtain multi-color photometry and proper motion of single stars in the Local Group. At the same time, slitless spectra can also provide effective temperature, metal abundance, surface gravity acceleration, low-precision radial velocity and other information for some stars ($r<23$ mag). CSST wide wavelength coverage ($255-1000$ nm) can enable observations to cover more star types and study the basic parameters of stars more comprehensively.
@@ 6.4.1 Star formation
Using CSST-TS, we can conduct THz spectrum mapping observations on different samples, including nearby molecular clouds, objects at different evolutionary stages, the central molecular belt of the Milky Way, and M31. We can obtain 492 GHz neutral carbon spectrum data and some HDO spectrum data. We also can perform gas circulation research in the Milky Way and M31 by combining spectral data from other observation equipment (such as CO, CII, HI, radio recombination lines, etc.), radiation transfer models, and astrochemical evolution simulations.
@@
Combined with mature accretion models, the accretion rate of young stars can be measured by using CSST UV data. The disk wind can be detected by observing the distribution of [OI]$\lambda$6300 using the CSST-IFS, and the generation mechanism of the disk wind will be analyzed by comparison with the relevant model [114, 115].
@@ 6.4.2 Massive stars and late evolution
For the pre-main sequence evolution of massive stars, young embedded clusters in the Galactic disk provide good observational samples. The angular diameter of these embedded clusters is comparable to the FoV of MCI. The high-resolution images of MCI can resolve each member star in these embedded clusters. Even in the case of severe extinction, MCI can detect massive young stars with an extinction of 20 mag within a range of 3 kpc. The wide spectral coverage of the IFS ($0.35-1.0$ $\mu$m) is conducive to the precise measurement of the spectral type, extinction and luminosity of massive young stars. Using light variation data from multiple telescopes, such as CSST, Kepler, and the Transiting Exoplanet Survey Satellite (TESS), combined with $Gaia$ and LAMOST data, we can study the internal structure and evolution of massive stars through asteroseismological methods, especially to advance our understanding of the laws of internal rotation and angular momentum transfer of stars.
@@ 6.4.3 Binary star evolution and special objects
CSST’s wide wavelength range ($255-1000$ nm), especially the coverage of the UV band, can cover binaries with larger masses and mass ratios, and realize accurate measurement of the parameters of the primary and companion stars. The magnitude limit of about 26 mag enables the sample to cover at least LMC/SMC (100 kpc), and is expected to cover M31/M33 (1Mpc), and greatly enrich the binary star sample of $100010<a<$ 1000 AU. Resolvable binaries can be obtained through CSST’s high-resolution imaging and milliarcsecond-level astrometric information (proper motion and parallax). The design of multi-color imaging and slitless spectroscopic survey arrangement can ensure that the binary star sample has good intrinsic homogeneity. CSST can construct a binary star sample with the largest number, good consistency, and high-precision stellar physical parameters.
@@ 6.5 Exoplanets
Exoplanets are a frontier and important field in astronomy. From the early 1990s to the early 21th century, the radial velocity method was mainly used to detect exoplanets. Since the 21st century, the transit method (such as Kepler, TESS, etc.) has become the main means of discovering exoplanets. Combined with microlensing, direct imaging, astrometry and other methods, a number of medium- and long-period exoplanets have been discovered, improving the completeness of medium- and long-period exoplanet samples. At present, more than 5,000 exoplanets have been discovered, and a relatively rich sample of exoplanets has been accumulated. In addition, combined with the spectral detection of the atmosphere of exoplanets, the spectral observation of the atmospheric components and contents are carried out to help us understand the evolution of planetary atmospheres. Existing and future space telescopes will further expand the planetary samples, improve the completeness of detection, characterize the atmospheric characteristics of exoplanets, and help us understand the diversity of planetary systems and the universal laws of formation and evolution.
@@ 6.5.1 Direct imaging detection
Direct imaging technology can truly “see” planets by directly observing planetary radiation, providing the possibility for studying characteristic signals of planetary atmospheres, and is a key technology for confirming signals of extrasolar life. So far, ground-based exoplanet imagers have discovered a total of about 100 exoplanets, most of which are more than 10 times the mass of Jupiter. Most of the planetary systems are very young ($1-200$ Myr) and have effective temperatures close to or exceeding 1000 K. These planets also have very strong radiation, which makes the imaging contrast requirements relatively low ($10^{-4}\sim 10^{-6}$ for infrared band). At present, ground-based observations cannot detect mature, low-temperature “cold” exoplanet.

### v3

@@ 6.2.3 Low-redshift galaxies
The CSST main survey is expected to measure the morphology and structure properties of young globular clusters of different galaxy types and environments within hundreds of Mpc. The high-quality imaging capability of the CSST will provide an unprecedented large sample of galaxy nuclear star clusters (NSCs). The high-precision photometric capability will obtain the NSC color index, radius, mass, stellar population, star formation history, morphology, and complete distribution of NSC physical parameters [126]. In this way, the correlation with the parameters of AGN dust ring can be established, the physical connection between the NSCs, bulges and massive black holes in a galaxy can be explored, and the physical process from the galaxy to the AGN transition region can be revealed.
@@
The outer regions of disk galaxies contain unique information about the physical processes that affect galactic evolution. Using the CSST high spatial resolution images, we can detect star clusters and single bright stars in the outer disk of galaxies. Considering the large sky coverage of CSST, the data at different radii of galaxies can be stacked, so that the areas with low star surface density can be detected. This can be helpful to answer two questions: the extension scale of galaxies and the physical mechanism that causes the growth of stellar disks. In addition, many galaxies have extended ultraviolet disks (XUV disks) at their outer edges [127]. CSST can be used to conduct UV-optical observations of a large number of XUV galaxies that can resolve single stars and study their physical properties [128].
@@
Star formation in galaxies is one of the main driving forces of galaxy growth and evolution, and is closely related to the mass structure, nuclear activity and other physical characteristics of galaxies. CSST’s multi-color imaging survey provides near-UV high-resolution deep-field images, which can effectively detect and analyze small-scale (<1<1 kpc) star formation and distribution in low-$z$ galaxies. CSST multi-color imaging survey is also expected to quantify the relationship between the star formation rate and molecular gas density on the sub-kpc scale, by exploring the group behavior and statistical properties of star formation in low-$z$ galaxies. Besides, CSST can conduct analyses on galaxies with strong radiation in the UV spectrum, which will help explain their origins [129].
@@ 6.2.4 Medium and high-redshift galaxies
When the resolution and S/N of galaxy images are low (such as when the galaxy is close to the detection limit), there are large errors in the measurement of its structural parameters. Therefore, it is necessary to use simulated images to test how the measurement errors of galaxy structural parameters vary with resolution and S/N. The mock galaxy images can be obtained by using galaxy image analysis code (e.g. GALFIT) or real multi-band galaxy images (e.g. HST CANDELS) [134, 133]. By applying the measurement methods, we can obtain the trend of measurement deviations of different structural parameters of galaxies as the image resolution and S/N change, which will help to understand and correct the statistical analysis results of galaxy image structural parameters.
The morphology of a galaxy characterizes its internal structure and its evolutionary stage. Galaxies of different morphological types may have different formation and evolution processes. The CSST image survey will significantly increase the number of samples of medium- and high-$z$ galaxies, and study their formation and evolution history of the morphology. It is expected that the statistical distribution characteristics of galaxy morphology within the highest redshift of 2 can be constructed, and the overall evolution process of the Hubble sequence can be reproduced [135, 136, 137, 138]. Besides, galaxy morphological parameters have rich physical meanings, and they are related to the star formation history of galaxies, intergalactic interactions, and small-scale structures.
@@
CSST will study the formation and evolution of substructures of medium and high-$z$ galaxies [139, 140]. The multi-component decomposition and statistical analysis of medium- and high-$z$ galaxy images will be explored, and the deviation of multi-component decomposition can be evaluated by using CSST simulated image data. It is expect to obtain a multi-component decomposition catalog of a large sample of galaxy images, and conduct statistical analysis on physical parameters, such as structural parameters of bulges and disks, bulge mass fraction, and stellar population in galaxies of different masses and star formation rates, to reveal their formation and evolution history as a function of redshift. In addition, CSST will focus on the identification and measurement of bars and spiral arms in galaxies within $z=1$. By combining HST and JWST data, CSST can identify and measure the bar structure in the early stages of disk galaxies and reveal the role of substructures in galaxies in their mass growth and structural evolution. CSST will also study the physical properties and evolution of clumps within galaxies.
@@ 6.2.5 Galaxy evolution and assembly history
Using CSST multi-band high-resolution photometric and slitless spectroscopic surveys, combined with other multi-band survey data, we are able to measure important parameters, such as stellar mass, star formation rate, dust extinction, stellar population age, and metallicity, for a large number of galaxies at the low and medium redshifts [141]. CSST can also measure the overall properties of a large number of complete galaxy samples, combined with galaxy pairs, merging galaxy statistics, and different local environment measurements, to construct the mass/luminosity function at different redshifts, as well as the scaling relationship between the mass-structure parameters-stellar population properties. Combined with numerical simulations or different types of galaxy evolution models, statistical modeling of these observations will greatly deepen our understanding of the galaxy stellar assembly history.
@@
Besides, since CSST can perform high-resolution wide-field survey, it will effectively probe the morphology and structure of galaxies and their dependence on the environment. CSST can carry out measurements of environments at different scales (including groups, filaments, voids, etc.), and study the impact of the environment on various properties of galaxies (including mass function, star formation rate, morphology, halo assembly history, gas, etc.), and their redshift evolution [142, 143, 144, 145]. CSST can also use its high-resolution photometric and spectroscopic survey to explore the formation and evolution history of galaxy groups and clusters, the formation and assembly history of the member galaxies, and the environmental effects of different redshifts and different types of galaxy groups/clusters on the formation and evolution of their member galaxies.
@@ 6.3.1 Stellar population
As a galaxy in the nearby Universe, the study of stellar populations in the Milky Way has made great progress in recent years. The release of the $Gaia$ astrometric catalog combined with ground-based stellar spectral survey data, such as the Large Sky Area Multi-Object Fiber Spectroscopic Telescope (LAMOST) [150] and SDSS [3], provides important observational data for the study of the Milky Way. The study of stellar populations in the Milky Way is mainly carried out in star clusters, field stars, nuclear bulges, etc.
Some studies indicate that the Milky Way may be special in the properties of star populations than other galaxies [151], so the understanding of the formation and evolution history of galaxies cannot rely solely on the results of the Milky Way. The statistical results of star populations in different types of galaxies will be helpful to have a deeper understanding of the specific process of galaxy formation and evolution [152].
@@
Based on this sample, great progress will be made in the properties of various stellar populations in nearby galaxies. Important observational results of the variation of the initial mass function in different interstellar environments will be obtained [153]. The evolution of stellar populations in different galaxies and the characteristics and evolution of stellar halos in different types of galaxies can be accurately measured. The number and types of satellite galaxies, and the number, types and spatial distribution of star clusters around different types of galaxies will be probed. The fundamentals of satellite galaxies and star clusters of different types of galaxies will be established, providing unprecedented complete data for the origin, merger and accretion process of satellite galaxies and constraints on the formation and evolution of star clusters.
@@ 6.3.2 Extinction and dust
The dust of various sizes, ranging from sub-nanometer to micron, is distributed in various environments from Solar System to the early Universe, and plays an important role in astrophysics research [154]. Although interstellar dust accounts for only one thousandth of the baryonic mass of the Milky Way, it interacts with electromagnetic radiation in almost the entire frequency domain, thus comprehensively affecting all branches of astrophysics. The absorption and scattering of starlight by dust is collectively called interstellar extinction. The measurement of interstellar extinction and infrared re-radiation of the Milky Way and extragalactic galaxies is the key to restore the intrinsic brightness and SED of astronomical sources and determine the star formation rate on a galaxy scale.
@@
Using the CSST multi-band photometric and slitless spectroscopic observations of the Milky Way and nearby galaxies in multiple sky regions and interstellar sight lines, combined with the infrared and millimeter data and interstellar element abundances from the Infrared Astronomical Satellite (IRAS), Spitzer and $Planck$, interstellar dust models can be constructed. The size distribution, chemical composition and other characteristics of dust can be obtained. The relationship between dust properties and the interstellar physical and chemical environment will be analyzed, and the relationship between dust extinction and other important interstellar observational features (such as interstellar diffuse belts, aromatic infrared radiation spectra, X-ray “halos”, etc.) can be studied. The electronic absorption spectral characteristics of carbon dust in the CSST UV band, such as graphene, carbon nanotubes, nanodiamonds, fullerenes (such as C60), carbon chains and non-crystalline carbon, will be examined [155, 156, 157]. In short, CSST will bring unprecedented opportunities for the study of extinction and dust physics.
@@ 6.4 Stars
CSST will obtain photometric data of billions of stars and hundreds of millions of stellar spectra during the ten-year survey. Its high spatial resolution ($0.074^{\prime\prime}$/pixel) and deep survey depth ($r\simeq 26$ mag) will enable us to observe single stars in the Milky Way and dozens or hundreds of nearby galaxies, such as the Andromeda Galaxy and the Triangulum Galaxy. By utilizing the unique advantages of CSST in the near-IR and near-UV bands and combining with other survey projects (such as LSST, HST, $Gaia$, etc.), we can obtain multi-color photometry and proper motion of single stars in the Local Group. At the same time, slitless spectra can also provide effective temperature, metal abundance, surface gravity acceleration, low-precision radial velocity and other information for some stars ($r<23$ mag). CSST wide wavelength coverage (2551000255-1000 nm) can enable observations to cover more star types and study the basic parameters of stars more comprehensively.
@@ 6.4.1 Star formation
Using CSST-TS, we can conduct THz spectrum mapping observations on different samples, including nearby molecular clouds, objects at different evolutionary stages, the central molecular belt of the Milky Way, and M31. We can obtain 492 GHz fine structure line of neutral carbon and rotational transitions of HDO in the same band. We will study gas circulation in the Milky Way and M31 by combining spectral data (such as CO, CII, HI, radio recombination lines, etc.) from other instruments, radiation transfer models, and astrochemical simulations.
@@
[…]
@@ 6.4.2 Massive stars and late evolution
[…]
@@ 6.4.3 Binary star evolution and special objects
CSST’s wide wavelength range (2551000255-1000 nm), especially the coverage of the UV band, can cover binaries with larger masses and mass ratios, and realize accurate measurement of the parameters of the primary and companion stars. The magnitude limit of about 26 mag enables the sample to cover at least LMC/SMC (100 kpc), and is expected to cover M31/M33 (1Mpc), and greatly enrich the binary star sample of $10<a<1000$ AU. Resolvable binaries can be obtained through CSST’s high-resolution imaging and milliarcsecond-level astrometric information (proper motion and parallax). The design of multi-color imaging and slitless spectroscopic survey arrangement can ensure that the binary star sample has good intrinsic homogeneity. CSST can construct a binary star sample with the largest number, good consistency, and high-precision stellar physical parameters.
@@ 6.5 Exoplanets
Exoplanets are a frontier and important field in astronomy. From the early 1990s to the early 21th century, the radial velocity method was mainly used to detect exoplanets. Since the 21st century, the transit method (such as Kepler, TESS, etc.) has become the main means of discovering exoplanets [175, 166]. Combined with microlensing, direct imaging, astrometry and other methods, a number of medium- and long-period exoplanets have been discovered, improving the completeness of medium- and long-period exoplanet samples. At present, more than 5,000 exoplanets have been discovered, and a relatively rich sample of exoplanets has been accumulated. In addition, combined with the spectral detection of the atmosphere of exoplanets, the spectral observation of the atmospheric components and contents are carried out to help us understand the evolution of planetary atmospheres. Existing and future space telescopes will further expand the planetary samples, improve the completeness of detection, characterize the atmospheric characteristics of exoplanets, and help us understand the diversity of planetary systems and the universal laws of formation and evolution.
@@ 6.5.1 Direct imaging detection
Direct imaging technology can truly “see” planets by directly observing planetary radiation, providing the possibility for studying characteristic signals of planetary atmospheres, and is a key technology for confirming signals of extrasolar life. So far, ground-based exoplanet imagers have discovered a total of about 100 exoplanets, most of which are more than 10 times the mass of Jupiter. Most of the planetary systems are very young (12001-200 Myr) and have effective temperatures close to or exceeding 1000 K. These planets also have very strong radiation, which makes the imaging contrast requirements relatively low ($10^{-4}\sim 10^{-6}$ for infrared band). At present, ground-based observations cannot detect mature, low-temperature “cold” exoplanet.